eCourses Fall2018_MGMT5123201-1910-10776 Question 3 A project with 7 activities is shown below Not yet Marked out of 1.00 F Flag FIN The duration of activities A B, C, and D are 20, 20, 15, and 27 days, respectively On the other hand, the duration of activities E, F, and G are 18, 21, and 32 days Assuming the project starts at the end of day o, the project completion time will be days Previous page
Answer #1

The path and duration are:

ACF = 20+15+21 = 56

BDF = 20+27+21 = 68

BDG = 20+27+32 = 79

BEG = 20+18+32 = 70

Project completion time will be highest path time = 79